OpenStudy (anonymous):

medal -Do you know the answer for this one .. What is the value of z so that -9 and 9 are both solutions of x2+z =103 ? a.-22 I think it is A b.3 c.22 d.184

OpenStudy (anonymous):

yea its a

OpenStudy (anonymous):

x 2 =103−z

OpenStudy (anonymous):

@blueweek77 what do u think a ?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

I think your right

OpenStudy (luigi0210):

-22? Well since you have a squared function, it will always be positive no matter the value \(( \pm9)^2-22=103\) So \(81-22=103\), does that seem right?

OpenStudy (anonymous):

Ok I see what you mean so it's just plain 22

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Yup~

OpenStudy (anonymous):

TY